OpenStudy (anonymous):

write an equation in point-slope form of the line that passes through the given points (7,2), (2,12)?

OpenStudy (mertsj):

slope = (12-2)/(2-7)=10/-5=-2 Now plug into point-slope form. Choose either point: y-2=-2(x-7) y-2=-2x-14 y=-2x-12